comet has an aphelion distance of 6 .1 ×10¹¹ m and periphelion distance of 5.1× 10¹¹m. The mass of the sun is 2.0 ×10^30 calculate the speed of the comet at the perihelion and the aphelion.
1
Expert's answer
2020-02-24T10:57:10-0500

1) The speed of the comet at the perihelion 


"v^2_p=GM\left(\frac{2}{r_p}-\frac{1}{a}\right)"

"r_p=a(1-e)"

"\frac{r_p}{r_a}=\frac{5.1}{6.1}=\frac{(1-e)}{(1+e)}\to e=0.0893"

"5.1\cdot 10^{11}=a(1-0.0893)"

"a=5.6\cdot 10^{11}\ m"

"v^2_p=(6.67\cdot 10^{-11})(2\cdot 10^{30})\left(\frac{2}{5.1\cdot 10^{11}}-\frac{1}{5.6\cdot 10^{11}}\right)"

"v_p=1.7\cdot 10^{4}\frac{m}{s}"

2) The speed of the comet at the aphelion


"v^2_a=GM\left(\frac{2}{r_a}-\frac{1}{a}\right)"

"v^2_p=(6.67\cdot 10^{-11})(2\cdot 10^{30})\left(\frac{2}{6.1\cdot 10^{11}}-\frac{1}{5.6\cdot 10^{11}}\right)"

"v_a=1.4\cdot 10^{4}\frac{m}{s}"